# Towards Reliable Connectivity: Measurement-Driven Assessment of Starlink and OneWeb Non-Terrestrial and 5G Terrestrial Networks

### v1

@@ Abstract
The emergence of commercial satellite communications networks, such as Starlink and OneWeb, has significantly transformed the communications landscape over the last years. As a complement to terrestrial cellular networks, non-terrestrial systems enable coverage extension and reliability enhancement beyond the limits of conventional infrastructure. Currently, the high reliance on terrestrial networks exposes communications to vulnerabilities in the event of terrestrial infrastructure failures, e.g., due to natural disasters. Therefore, this work proposes the joint evaluation of Key Performance Indicators (KPIs) for two non-terrestrial satellite networks (Starlink and OneWeb) and two terrestrial cellular networks to assess the current performance of these technologies across three different environments: (i) urban, (ii) suburban, and (iii) forest scenarios. Additionally, multi-connectivity techniques are explored to determine the benefits in connectivity when two technologies are used simultaneously. For instance, the outage probability of Starlink and OneWeb in urban areas is reduced from approximately 12-21% to 2% when both solutions are employed together. Finally, the joint analysis of KPIs in both terrestrial and non-terrestrial networks demonstrates that their integration enhances coverage, improves performance, and increases reliability, highlighting the benefits of combining satellite and terrestrial systems in the analyzed environments.
publicationid: pubid:
@@ I Introduction
Given the respective benefits and costs of each communications system, the state-of-the-art includes multiple studies that quantify the performance and the quality of TN and NTN under different assumptions, using simulations as well as real-world measurements. Focusing on terrestrial networks, there is extensive work related to the simulation of KPIs in the radio and network layers for 5G networks. For instance, Kang et al. present an extensive simulation analysis of the feasibility of the upper mid-band for use in cellular networks [24]. Iranmanesh et al. study capacity improvement in 5G networks through mobile base stations located on drones [25]. Vielhaus et al. propose a system-level simulator, which allows realistic network deployments to be reproduced in simulation [26]. Both theoretical advances and the development of simulators are fundamental to the practical deployment of cellular networks. From an experimental point of view, studies such as those by Aerts et al. [27] and Chiaraviglio et al. [28] conduct empirical studies of the physical layer to determine the levels of electromagnetic fields generated by 5G deployments. Additionally, Mohebi et al. [29] report network performance in terms of latency and throughput for commercial 5G networks. Concerning non-terrestrial networks, multiple efforts have been made in relation to stochastic theoretical modeling and simulation of satellite communication systems, such as the work of Jung et al. [30] and Abdelsadek et al. [31]. Focusing on experimental work, empirical performance evaluations of commercial global constellations have also recently begun to be explored, such as the examples in the work of Laniewski et al. [32, 33]. Note that due to the recent deployment of these networks, the experimental state-of-the-art related to NTN networks remains relatively scarce compared to both theoretical research in NTN and the extensive state-of-the-art on terrestrial networks. Taking a step further and combining the fields described above, we have conducted preliminary studies jointly analyzing the performance of cellular networks and Starlink in rural areas [34] and high-mobility environments along highways [35].
@@
In this paper, we propose a joint empirical analysis of the performance of up to four commercial communications systems around the city of Copenhagen (Denmark): two major Danish MNOs, and two NTN satellite systems, Starlink and OneWeb. Thus, the aim is to assess the feasibility of their deployment across multiple environments, such as urban, suburban/residential, and rural/forest environments. The evaluation focuses on Key Performance Parameters (KPIs) related to data rate and latency in both directions, downlink and uplink. While optimal performance of the terrestrial network is expected in such urban scenarios, the study shows a comparison between well-established terrestrial systems and emerging non-terrestrial alternatives. In addition, the study provides an estimate of the expected performance if multiple combinations of the aforementioned systems were integrated. Finally, the assessment focuses both on mobility conditions, where the communications system is located in a vehicle that is constantly in motion, and on static conditions, where the equipment is fixed in a specific location. Overall, the main contributions of this work are:
@@
  • Development of a mobility measurement campaign for the joint evaluation of TN 5G, and NTN Starlink and Oneweb commercial networks in urban, suburban, and forest scenarios.
@@
  • Statistical analysis of KPIs that impact end-user performance, such as latency and throughput in both uplink and downlink.
@@
  • Evaluation of network performance improvement in the case of integrating multiple technologies through multi-connectivity, e.g, Cell./Cell., Cell./Sat., or Sat./Sat. multi-connectivity.
  • Comparison of connectivity performance in commercial satellite networks for NLoS in static conditions.
@@ II Measurement Campaign and Scenario
This campaign simultaneously evaluates two cellular connectivity solutions, corresponding to Denmark’s two main 5G cellular operators, as well as the non-terrestrial networks Starlink and OneWeb. At the time of performing the measurement campaign, both mobile network operators have a cellular network of hundreds of base stations within the region under study [36]. With regard to satellite operators, Starlink has approximately 8000 satellites in LEO orbits between 550 km and 700 km [37], while OneWeb has approximately 648 satellites deployed in LEO orbits at an altitude of 1200 km [38]. In the case of Starlink, the vast majority of satellites are deployed in orbits with an inclination of 53º since the largest proportion of the global population is located below these latitudes. In addition, new polar orbits have been deployed to provide coverage at higher latitudes. In the case of the city of Copenhagen, located at a latitude of 55º, this deployment is sufficient to ensure visibility with multiple satellites simultaneously. In the case of OneWeb, the constellation orbits exclusively in polar orbits around 87º, which means better coverage at high latitudes compared to the equator. For the latitude at which the study area is located, this deployment is sufficient to ensure continuous coverage if there are no obstructions on the horizon.
Focusing on the physical characteristics of each scenario, Figs. 2(a)-(c) show representative photographs of the environments corresponding to various locations along the routes covered. The urban environment is characterized by avenues and streets lined with buildings of four to five stories high. Due to the high population density in this area, MNOs have a highly densified network of base stations. According to statistics from the Danish Agency for Digital Government, there are approximately 15 base stations per km2 in this area, indicating a high density of cell sites [36]. In the context of Starlink and OneWeb constellations, although the coverage is global, this scenario is prone to signal blockage at low elevation angles due to the presence of buildings, requiring the satellite to establish connectivity in regions close to the zenith. On the other hand, the suburban scenario is characterized by a landscape consisting of single-family homes and one to two-story buildings. Due to the lower population density compared to urban scenarios and more favorable radio propagation conditions, the number of base stations in this area drops to 3-5 base stations per km2 [36]. However, the landscape is more favorable for satellite constellations given the lower number of obstacles on the horizon, increasing the probability of LoS between the ground equipment and the satellite. Finally, the forest scenario is characterized by high vegetation density even at the zenith, which implies challenging conditions for radio propagation in satellite systems. Similarly, the area under study is located approximately one kilometer from the closest population center, with only one base station from each operator within a two-kilometer radius [36].
@@
III Measurement Setup
@@
For the simultaneous evaluation of the two cellular networks and the two satellite networks, a dedicated setup is required. In particular, given the mobility conditions of the measurement campaign, the user equipment needs to be mounted on a vehicle. Specifically, two vehicles are used due to space limitations. The first one is equipped with the user equipment for the cellular and Starlink networks, and the second one is equipped with the user equipment for OneWeb. Fig. 3 shows the user equipment for the four networks involved, mounted on the two vehicles driven. During the measurement campaign, both vehicles were driven together at a short distance to preserve similar physical conditions among the four connectivity interfaces. Regarding the cellular connectivity framework, the experimental setup comprises two Intel NUC (Next Unit of Computing) devices. Each device is equipped with two onboard Gigabit Ethernet interfaces and expanded with two USB 3.1 Gigabit Ethernet adapters. These NUCs operate as traffic generators, hosting a UDP client that synchronizes with a remote UDP server located at the AAU premises. To establish connectivity, two Teltonika RUTX50 Industrial 5G routers were deployed to provide high-performance and independent network interfaces to one NUC. Routers are configured to prioritize the operator’s 5G network, automatically falling back to 4G when 5G coverage is not available. Each router connects to four omnidirectional antennas with 5 dBi gain (model AOA-4G-5W-SMM) depicted in Fig. 3. As for the Starlink setup, it employs a Starlink Standard Actuated Gen 2 antenna (51.3 $\times$ 30.3 cm). A Starlink Gen 2 router serves as the gateway, establishing IP connectivity and linking the satellite backhaul to the NUC via an Ethernet interface. Regarding the OneWeb setup, a Kymeta Hawk U8 terminal (89.5 $\times$ 89.5 cm) is employed. This device features a built-in Ethernet interface that connects directly to the OneWeb backhaul, thereby providing the NUC with satellite internet access. To ensure precise time alignment, a GPS-based NTP source distributes a Pulse Per Second (PPS)-driven time reference via Ethernet for each NUC, thus enabling microsecond accuracy time synchronization. In addition to all of the above, each vehicle is equipped with batteries and DC/AC inverters power electronic components, laptops to monitor measurements in real time, and GPS antennas to keep track of the location.
@@
Regarding the measurement procedures, this setup evaluates four Key Performance Indicators: (i) downlink one-way delay (OWD DL), (ii) uplink one-way delay (OWD UL), (iii) downlink throughput (TP DL), and (iv) uplink throughput (TP UL). For this purpose, UDP packets are transmitted in both directions at 2 ms intervals, thus measuring the one-way delay as the difference between the time of sending and receiving on the server and client (DL), respectively, and vice versa (UL). This short interval between packets ensures constant monitoring of the link. Additionally, the UDP packet size is adjusted so that the traffic corresponds to 5/5 Mbps DL/UL, thus establishing a target data rate of 5 Mbps in both directions. This target data rate has been chosen since it is the throughput required in the CFB for high-resolution real-time video transmission in emergency situations. In short, this setup allows the measurement of 4 KPIs/interface for the two cellular operators, Starlink and OneWeb, thus enabling the simultaneous and synchronized measurement of a total of 16 KPIs.

### v2

@@ Abstract
The emergence of commercial satellite communications networks, such as Starlink and OneWeb, has significantly transformed the communications landscape over the last years. As a complement to terrestrial cellular networks, non-terrestrial systems enable coverage extension and reliability enhancement beyond the limits of conventional infrastructure. Currently, the high reliance on terrestrial networks exposes communications to vulnerabilities in the event of terrestrial infrastructure failures, e.g., due to natural disasters. Therefore, this work proposes the joint evaluation of Key Performance Indicators (KPIs) for two non-terrestrial satellite networks (Starlink and OneWeb) and two terrestrial cellular networks to assess the current performance of these technologies across three different environments: (i) urban, (ii) suburban, and (iii) forest scenarios. Additionally, multi-connectivity techniques are explored to determine the benefits in connectivity when two technologies are used simultaneously. For instance, the outage probability of Starlink and OneWeb in urban areas is reduced from approximately 12-21% to 2% when both solutions are employed together. Finally, the joint analysis of KPIs in both terrestrial and non-terrestrial networks demonstrates that their integration enhances coverage, improves performance, and increases reliability, which highlights the benefits of combining satellite and terrestrial systems in the analyzed environments.
@@ I Introduction
Given the respective benefits and costs of each communications system, the state-of-the-art includes multiple studies that quantify the performance and the quality of TN and NTN under different assumptions, using simulations as well as real-world measurements. Focusing on terrestrial networks, there is extensive work related to the simulation of KPIs in the radio and network layers for 5G networks. For instance, Kang et al. present an extensive simulation analysis of the feasibility of the upper mid-band for use in cellular networks [24]. Iranmanesh et al. study capacity improvement in 5G networks through mobile base stations located on drones [25]. Vielhaus et al. propose a system-level simulator, which allows realistic network deployments to be reproduced in simulation [26]. Both theoretical advances and the development of simulators are fundamental to the practical deployment of cellular networks. From an experimental point of view, studies such as those by Aerts et al. [27] and Chiaraviglio et al. [28] conduct empirical studies of the physical layer to determine the levels of electromagnetic fields generated by 5G deployments. Additionally, Mohebi et al. [29] report network performance in terms of latency and throughput for commercial 5G networks. Concerning non-terrestrial networks, multiple efforts have been made in relation to stochastic theoretical modeling and simulation of satellite communication systems, such as the work of Jung et al. [30] and Abdelsadek et al. [31]. Focusing on experimental work, empirical performance evaluations of commercial global constellations have also recently begun to be explored, such as the examples in the work of Laniewski et al. [32, 33]. In addition, Pan et al. [34, 35] have thoroughly characterized Starlink by analyzing its routing behavior, network architecture, and backbone infrastructure, as well as performance metrics acquired from terminals distributed worldwide. Similarly, Zhao et al. [36] provide a cross-layer analysis of OneWeb ground architecture and the impact of traffic routing through the satellite network on network KPIs. Recent state-of-the-art has also evaluated the performance of NTN networks under different traffic patterns, focusing on the impact of transport-layer protocols. For instance, Careau et al. [37] have performed a comparative analysis of TCP and UDP traffic over the Starlink network. Similarly, García et al. [38] have studied several congestion control variants over Starlink, showing that protocol selection can strongly affect network performance. Note that due to the recent deployment of these networks, the experimental state-of-the-art related to NTN networks remains relatively scarce compared to the extensive state-of-the-art on terrestrial networks. Taking a step further and combining the fields described above, we have conducted preliminary studies jointly analyzing the performance of cellular networks and Starlink in rural areas [39] and high-mobility environments along highways [40].
@@
In this paper, we propose a joint empirical analysis of the performance of up to four commercial communications systems around the city of Copenhagen (Denmark): two major Danish MNOs, and two NTN satellite systems, Starlink and OneWeb. Thus, the aim is to assess the feasibility of their deployment across multiple environments that are prone to exhibit heterogeneous and potentially challenging radio propagation conditions due to radio obstruction, NLoS conditions, and limited satellite visibility, such as urban, suburban/residential, and rural/forest environments. The evaluation focuses on Key Performance Parameters (KPIs) related to data rate and latency in both directions, downlink and uplink. This unidirectional measurement approach enables the independent characterization of uplink and downlink paths, revealing latency and performance asymmetries that cannot be captured by conventional round-trip measurements. While optimal performance of the terrestrial network is expected in such urban scenarios, the study shows a comparison between well-established terrestrial systems and emerging non-terrestrial alternatives. In addition, the study provides an estimate of the expected performance if multiple combinations of the aforementioned systems were integrated, covering TN–TN, TN–NTN, and NTN–NTN multi-connectivity configurations. For instance, the analysis of exclusively satellite multi-connectivity enables the assessment of reliability gains achievable through satellite diversity. Finally, the assessment focuses both on mobility conditions, where the communications system is located in a vehicle that is constantly in motion, and on static conditions, where the equipment is fixed in a specific location. Overall, the main contributions of this work are:
@@
  • Development of a mobility measurement campaign for the joint evaluation of TN 5G, and NTN Starlink and Oneweb commercial networks in urban, suburban, and forest scenarios under challenging radio propagation conditions due to radio obstruction, NLoS conditions, and limited satellite visibility.
@@
  • Unidirectional statistical analysis of end-to-end KPIs that impact end-user performance, such as throughput and one-way delay latency metrics in both uplink and downlink, enabling the identification of link performance asymmetries that cannot be captured through round-trip measurements.
@@
  • Evaluation of network performance improvement in the case of integrating multiple technologies through multi-connectivity, including terrestrial-only (TN–TN), hybrid terrestrial and non-terrestrial (TN–NTN), and exclusively non-terrestrial (NTN–NTN) multi-connectivity configurations.
  • Assessment of the robustness of commercial satellite connectivity under NLoS conditions in both mobile and static scenarios, representative of challenging real-world deployment scenarios.
@@ II Measurement Campaign and Scenario
This campaign simultaneously evaluates two cellular connectivity solutions, corresponding to Denmark’s two main 5G cellular operators, as well as the non-terrestrial networks Starlink and OneWeb. At the time of performing the measurement campaign, both mobile network operators have a cellular network of hundreds of base stations within the region under study [41]. With regard to satellite operators, Starlink has approximately 8000 satellites in LEO orbits at the time of measurement between 550 km and 700 km [42], while OneWeb has approximately 648 satellites deployed in LEO orbits at an altitude of 1200 km [43]. Regarding inter-satellite links (ISLs), Starlink has introduced ISL capabilities in its newer generations of satellites, while OneWeb does not currently support ISLs. In the case of Starlink, the vast majority of satellites are deployed in orbits with an inclination of 53º since the largest proportion of the global population is located below these latitudes. In addition, new polar orbits have been deployed to provide coverage at higher latitudes. In the case of the city of Copenhagen, located at a latitude of 55º, this deployment is sufficient to ensure visibility with multiple satellites simultaneously. In the case of OneWeb, the constellation orbits exclusively in polar orbits around 87º, which means better coverage at high latitudes compared to the equator. For the latitude at which the study area is located, this deployment is sufficient to ensure continuous coverage if there are no obstructions on the horizon. Note that the results shown throughout the work are specific to the 55º Copenhagen high latitude scenario, and although both constellations provide coverage at this latitude, their relative performance might differ in terms of the latitude. In particular, OneWeb’s relative performance is expected to improve with increasing latitude due to its near-polar orbital configuration, while Starlink benefits from higher satellite density at lower latitudes closer to the equator. Regarding the network topology, the scientific community has carried out extensive work in mapping the terrestrial infrastructure for NTN networks, i.e., the Ground Stations or Satellite Network Portals (SNPs) and Points of Presence (PoPs). For instance, the works of Wang et al. [44] and Zhao et al. [36] have identified this infrastructure for Starlink and OneWeb, respectively. Considering our geographical position in Copenhagen, the Starlink ground stations in Aerzen (Germany), Frankfurt (Germany) and Hoofddorp (Netherlands), and the Starlink PoP in Frankfurt (Germany) are noteworthy due to their geographical proximity. For OneWeb, the nearest ground stations are geographically sparser in Sintra (Portugal), Palermo (Italy), Stara Zagora (Bulgaria) and Piteå (Sweden), while the PoP closest to our location is in Amsterdam (Netherlands). Therefore, the larger distance between the ground infrastructure and our location indicates higher expected end-to-end latency in OneWeb compared to Starlink. Note also that, from a global perspective, the development of ground infrastructure is more mature for Starlink, with 49 PoPs compared to OneWeb’s 29 PoPs, both as of 2025 [44, 36]. Hence, it is more likely that a Starlink PoP is geographically closer to the user’s location given a random location. In terms of signal structure, Starlink uses Orthogonal Frequency Division Multiplexing (OFDM) [45], while OneWeb implements MultiFrequency Time Division Multiple Access (MF-TDMA) Single Carrier (SC) [46].
@@
Focusing on the physical characteristics of each scenario, Figs. 2(a)-(c) show representative photographs of the environments corresponding to various locations along the routes covered. The urban environment is characterized by avenues and streets lined with buildings of four to five stories high. Due to the high population density in this area, MNOs have a highly densified network of base stations. According to statistics from the Danish Agency for Digital Government, there are approximately 15 base stations per km2 in this area, indicating a high density of cell sites [41]. In the context of Starlink and OneWeb constellations, although the coverage is global, this scenario is prone to signal blockage at low elevation angles due to the presence of buildings, requiring the satellite to establish connectivity in regions close to the zenith. On the other hand, the suburban scenario is characterized by a landscape consisting of single-family homes and one to two-story buildings. Due to the lower population density compared to urban scenarios and more favorable radio propagation conditions, the number of base stations in this area drops to 3-5 base stations per km2 [41]. However, the landscape is more favorable for satellite constellations given the lower number of obstacles on the horizon, increasing the probability of LoS between the ground equipment and the satellite. The forest scenario corresponds to a narrow road surrounded by dense vegetation, where tall trees are found in both sides of the route. Tree crowns, branches, and leaves introduce significant obstruction of the sky, resulting in a predominantly blocked zenith and limited satellite visibility. In most sections, the dense vegetation causes strong attenuation and frequent blockage of satellite links, particularly at low and medium elevation angles. Limited visibility is occasionally available through small gaps in the foliage. In some short segments, partial visibility at very high elevation angles ($70\degree$) is observed. However, these conditions are transient and typically persist only for a few seconds due to mobility. Regarding the cellular network conditions, the area under study is located approximately one kilometer from the closest population center, with only one base station from each operator within a two-kilometer radius [41].
@@ III Measurement Setup
For the simultaneous evaluation of the two cellular networks and the two satellite networks, a dedicated setup is required. In particular, given the mobility conditions of the measurement campaign, the user equipment needs to be mounted on a vehicle. Specifically, two vehicles are used due to space limitations. The first one is equipped with the user equipment for the cellular and Starlink networks, and the second one is equipped with the user equipment for OneWeb. Fig. 3 shows the user equipment for the four networks involved, mounted on the two vehicles driven. During the measurement campaign, both vehicles were driven together at a short distance to preserve similar physical conditions among the four connectivity interfaces. Regarding the cellular connectivity framework, the experimental setup comprises two Intel NUC (Next Unit of Computing) devices. Each device is equipped with two onboard Gigabit Ethernet interfaces and expanded with two USB 3.1 Gigabit Ethernet adapters. These NUCs operate as traffic generators, hosting a UDP client that synchronizes with a remote UDP server located at the AAU premises. To establish connectivity, two Teltonika RUTX50 Industrial 5G routers were deployed to provide high-performance and independent network interfaces to one NUC. Routers are configured to prioritize the operator’s 5G network, automatically falling back to 4G when 5G coverage is not available. Each router connects to four omnidirectional antennas with 5 dBi gain (model AOA-4G-5W-SMM) depicted in Fig. 3. As for the Starlink setup, it employs a Starlink Standard Actuated Gen 2 antenna (51.3 $\times$ 30.3 cm). Due to mobility during the measurement campaign, this antenna was configured to maintain a fixed orientation toward the sky, avoiding mechanical reconfiguration. A Starlink Gen 2 router serves as the gateway, establishing IP connectivity and linking the satellite backhaul to the NUC via an Ethernet interface. Regarding the OneWeb setup, a Kymeta Hawk U8 terminal (89.5 $\times$ 89.5 cm) is employed. This device features a built-in Ethernet interface that connects directly to the OneWeb backhaul, thereby providing the NUC with satellite internet access. Differences between the antennas arise from the distinct market segments targeted by the two satellite operators, with OneWeb user terminal focusing on enterprise and government levels, and Starlink user terminal being designed primarily for the consumer market. The effect of using these two different antennas is discussed in detail in the following sections. To ensure precise time alignment, a GPS-based NTP source distributes a Pulse Per Second (PPS)-driven time reference via Ethernet for each NUC, thus enabling microsecond accuracy time synchronization. In addition to all of the above, each vehicle is equipped with batteries and DC/AC inverters power electronic components, laptops to monitor measurements in real time, and GPS antennas to keep track of the location.
@@
Regarding the measurement procedures, this setup evaluates four Key Performance Indicators: (i) downlink one-way delay (OWD DL), (ii) uplink one-way delay (OWD UL), (iii) downlink throughput (TP DL), and (iv) uplink throughput (TP UL). For this purpose, UDP packets are transmitted in both directions at 2 ms intervals, thus measuring the one-way delay as the difference between the time of sending and receiving on the server and client (DL), respectively, and vice versa (UL). Therefore, the OWD evaluates the end-to-end delay experienced by the user in both uplink (OWD UL) and downlink (OWD DL), accounting for both the physical propagation delay and network-induced delays, including routing, queuing, and handover effects. The short interval between packets ensures constant monitoring of the link. Additionally, the UDP packet size is adjusted so that the traffic corresponds to 5/5 Mbps DL/UL, thus establishing a target data rate of 5 Mbps in both directions. This target data rate has been chosen since it is the throughput required in the CFB for high-resolution real-time video transmission in emergency situations. Therefore, this target is representative of a broad class of applications relying on continuous uplink or downlink video streaming. Note that these values are not intended to evaluate the networks under maximum capacity or heavy traffic conditions, but rather to assess the reliability of network connectivity across the three considered measurement scenarios shown in Section II. In short, this setup allows the measurement of 4 KPIs/interface for the two cellular operators, Starlink and OneWeb, thus enabling the simultaneous and synchronized measurement of a total of 16 KPIs.
@@ IV-A GPS Traces
Fig. 6: Time trace of a 90-degree turn at an intersection in the urban area for satellite operators OneWeb (
) and Starlink (
@@
Analyzing the latency values for Op. A and Op. B in Fig. 5(a), it can be seen that the OWD is stable with uniform values between 15 and 25 ms throughout the different locations. This fact is especially remarkable in the case of the downlink. For the uplink OWD, although the values remain stable, there are several instances where latency increases to values between 100 and 120 ms (see Op. A). In any case, the low latency values obtained for terrestrial networks are to be expected given the redundancy of base stations in this area, where there are multiple cells covering the user terminal that can function as backup even if the main cell is not operating correctly or is congested. In the case of the uplink, the link budget can play a fundamental role in the occasional increase in latency, since the transmission power in the user equipment is certainly more restricted than the transmission power in the base station.
Regarding the satellite operators, there are multiple coverage interruptions along the route where both uplink and downlink OWD exceed several seconds. Given the characteristics of the route, the buildings on the different avenues and streets act as sources of LoS obstruction, causing these interruptions. In particular, after analyzing the time traces of the measurement campaign for the urban environment, it was observed that the main coverage interruptions occur when the vehicle turns at an intersection. The turning of the vehicle causes a change in the position of the satellites on the horizon from the user terminal perspective, so that during turns at intersections or curves, LoS between the satellite and the ground equipment is momentarily obstructed. As an illustrative example, Fig. 6 shows the downlink OWD time trace corresponding to a 90-degree turn in the urban area. While prior to the heading change, both connectivity options maintain a stable latency below 100 ms, after initiating the turn and changing the orientation of our vehicle, connectivity is completely lost with OneWeb for more than 20 seconds and intermittently with Starlink at intervals of around 1 second. This suggests that the change in the vehicle’s orientation during the turn changes the spatial distribution of the satellites from the user equipment perspective, which eventually causes a momentary loss of visibility with some of the satellites, resulting in a loss of connectivity for several seconds. Although this Fig. 6 shows a specific example, this behavior has been observed repeatedly throughout the measurements conducted in the urban area. In some cases, OneWeb is the most affected constellation, as shown in Fig. 6, while in other cases, Starlink is the most affected. This lack of correlation between the instants at which the visibility of each constellation is impaired due to the independence of the orbits of the satellites in each constellation will be exploited to increase the reliability of satellite networks throughout Section V.
In the case of throughput analysis, similar behavior to that shown by latency is found. Cellular operators show stable values around the targeted 5/5 Mbps in both downlink and uplink due to the high density of available cells. In the case of satellite operators, constant connection drops are observed, with throughput instantly falling from 5 Mbps to 0 Mbps, suggesting a sudden blockage of the LoS. It is still significant that these drops tend to occur after turns at intersections.
IV-B One Way Delay Statistics